OpenStudy (anonymous):

Simplify. Assume all variables are nonzeros (8x^3y/2)^2 PLEASE HELP !

OpenStudy (anonymous):

\[\frac{8x^3y}{2}=4x^2y\] is start then square that one

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

i divided eight by two and got four

OpenStudy (anonymous):

Ooh Okay. Got it. So once i do that what do i do next?

OpenStudy (anonymous):

\[(4x^3y)^2=16x^6y^2\] is the last step

OpenStudy (anonymous):

Okay Thanks so MUCH!!